# Optimizing Case-Based Reasoning System for Functional Test Script Generation with Large Language Models

### v1

@@ Abstract.
Case-Based Reasoning, Large Language Model, Test Script Generation, Functional Testing, Reinforcement Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation ema; June 03–05,2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language generation
@@ 1. Introduction
Figure 2. The overall paradigm of the proposed CBR+Re4 for functional test script generation with LLMs: (a) The CBR system; (b) Reranking-based retrieval finetuning; (c) Reinforced reuse finetuning.
@@ 3.2. Reinforced Reuse Finetuning
In the CBR system, the Reuse step focuses on adapting solutions of past similar cases to the current task. While modern LLMs demonstrate strong instruction-following capabilities, they may still suffer from misalignment issues (Wen et al., 2024; Tan et al., 2024) when confronted with unseen tasks. Therefore, further finetuning of LLMs is necessary to ensure alignment with the desired behavior in our production scenarios. To this end, we propose a reinforced reuse finetuning method that incorporates both supervised finetuning (SFT) and reinforcement learning finetuning (RLFT), as shown in Figure 2(c).
@@ 4.1.4. Baselines
We compare CBR+Re4 with four baselines:
@@ 4.2.1. Overall Comparison
We present the experimental results of the offline evaluation metrics for both datasets in Table 1. We can observe that Naïve SFT performs the worst among all methods, even underperforming the simplest baseline, Naïve CBR. This is expected, as functional test script generation requires the LLM to invoke functions beyond its knowledge, which cannot be directly injected by SFT. Consequently, this leads to significant hallucination issues during inference. In contrast, CBR outperforms Naïve CBR, thanks to the strong foundational capabilities of LLMs. These capabilities are further enhanced by incorporating SFT, making CBR+SFT the strongest baseline. Among all the methods, the proposed CBR+Re4 achieves the highest FF1 score for both datasets, showing improvement of 3.30% and 6.80% over the best baseline, CBR+SFT, for DCN and SP, respectively. This validates the effectiveness of both the reranking based retrieval finetuning method and the reinforced reuse finetuning method.
@@ 4.3.1. Analyses for Reranking-based Retrieval Finetuning
As bge-m3 is widely recognized as a strong zero-shot retriever model, most of our early attempts to optimize the Retrieve step resulted in negative effects, primarily due to the absence of ground-truth labels. As such, we analyze the proposed reranking-based retrieval finetuning method via an ablation study. The performance gap of CBR+Re4 without retrieval finetuning is reported for both datasets in Figure 5(a). The results show that the ablation causes a performance drop of approximately 1% in terms of FF1 score on both datasets. This highlights the effectiveness of the proposed reranking-based retrieval finetuning method, which is also the only attempt of our team that yielded a positive impact on the retriever model.
@@ 4.4. Online Evaluation Results
We compare the previously deployed CBR+SFT system with two methods: (1) CBR+Re4 and (2) CBR+Re4 w/o Retain, an ablation variant of the CBR system where the Retain step is removed and the case bank remains fixed. It is worth noting that CBR+Re4 w/o Retain can be viewed as a variant of RAG, as it only reserves the Retrieve-Reuse process in CBR. Figure 7 depicts the average improvement in FF1 score as the number of online samples increases across two PDUs, DCN and SP. We observe that CBR+Re4 w/o Retain significantly underperforms CBR+Re4. This performance gap can be attributed to the dynamic nature of our scenario, where the software is continuously updated, and new modules are regularly introduced. Consequently, updating the case bank in an online manner is crucial for maintaining the effectiveness of test script generation. The Retain step in the CBR system addresses this need by enabling seamless integration of new information into the case bank. Furthermore, the proposed CBR+Re4 consistently outperforms the previously deployed CBR+SFT, achieving improvements in FF1 score of 2.75% in DCN and 10.08% in SP. These results validate the effectiveness of the proposed Re4 optimization method. The variation in improvement across the two PDUs can be attributed to the nature of the test scripts. In DCN, the scripts are much more function-intensive, meaning even minor improvements in FF1 score represent a significant advancement.
@@ Appendix A Algorithmic Details
In this section, we provide more implementation details of the proposed business metrics, and then provide the pseudo-code of the proposed method.
A.1. Implementation details for Business Metric
@@ Appendix B Why Does REINFORCE Work in Our RLFT Setting?
Now, we provide empirical evidence supporting the above claim. Since a smaller gradient variance corresponds to a smaller gradient norm, we follow prior work (Li et al., 2024) and plot the gradient norm of REINFORCE during the RLFT stage in Figure 9. The results reveal that the gradient norm of REINFORCE in our RLFT settings remains below 1 across both datasets, whereas this value typically fluctuates between 10 and 1000 in RLHF (Li et al., 2024). Thus, REINFORCE does not suffer from the high variance issue and achieves superior performance in a simple yet effective manner.
Figure 9. Gradient norm of REINFORCE in the RLFT stage.
@@ Appendix C Discussion of the Compared RLHF Algorithms
In this section, we present a detailed discussion with the state-of-the-art RLHF algorithms compared in the main body of the paper.
@@
where $T_{i}$ denotes the number of tokens in $i$-th generated script, $\hat{\rho}_{i,t}=\frac{\pi_{\theta}(\hat{y}_{i,t}|q,\mathcal{C}_{q},\hat{y}_{i,<t})}{\pi_{\theta_{\text{old}}}(\hat{y}_{i,t}|q,\mathcal{C}_{q},\hat{y}_{i,<t})}$ denotes the importance rate of the $t$-th token in $i$-th generated test script, $\hat{A}_{i,t}=\frac{r(\hat{y}_{i})-\text{mean}(\{r(\hat{y}_{1}),r(\hat{y}_{2}),...,r(\hat{y}_{K})\})}{\text{std}(\{r(\hat{y}_{1}),r(\hat{y}_{2}),...,r(\hat{y}_{K})\})}$ denotes the group-relative reward for advantage estimation, $\epsilon$ denotes the clipping parameter. GRPO achieves better performance than the proposed REINFORCE algorithm, benefiting from $K\times$ on-policy samples per query. Different from (Shao et al., 2024) that sets $K=64$, we set $K=4$ in our implementation to accommodate computational constraints, which brings 4×4$\times$4 × time and computational costs compared to REINFORCE.

### v2

@@ Abstract.
Case-Based Reasoning; Large Language Model; Test Script Generation; Functional Testing; Reinforcement Learning
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1454-2/25/08doi: 10.1145/3711896.3737254ccs: Computing methodologies Natural language generation
@@ 1. Introduction
Figure 2. The overall paradigm of the proposed CBR+Re4. (a) The CBR system, which follows a 4R cycle of Retrieve, Reuse, Revise and Retain; (b) Reranking-based retrieval finetuning, which identifies cases with both high semantic and script similarity as positive examples and applies contrastive learning for finetuning; (c) Reinforced reuse finetuning, which consists of supervised finetuning and function f1 score based reinforcement learning finetuning for LLMs.
@@ 3.2. Reinforced Reuse Finetuning
In the CBR system, the Reuse step focuses on adapting solutions of past similar cases to the current task. While modern LLMs demonstrate strong instruction-following capabilities, they may still suffer from misalignment issues (Wen et al., 2024; Tan et al., 2024; Yan et al., 2025) when confronted with unseen tasks. Therefore, further finetuning of LLMs is necessary to ensure alignment with the desired behavior in our production scenarios. To this end, we propose a reinforced reuse finetuning method that incorporates both supervised finetuning (SFT) and reinforcement learning finetuning (RLFT), as shown in Figure 2(c).
@@ 4.1.4. Baselines
To the best of our knowledge, this is the first paper that discusses the optimization of the full workflow in the CBR system with LLMs. Thus, there are not any other previous works suitable for comparison. To contextualize the performance of the proposed method, we carefully design four baselines:
@@ 4.2.1. Overall Comparison
We present the experimental results of the offline evaluation metrics for both datasets in Table 1. Note that we follow most LLM works to report the result of a single run due to the overly computational costs, which might be one limitation of this work. We can observe that Naïve SFT performs the worst among all methods, even underperforming the simplest baseline, Naïve CBR. This is expected, as functional test script generation requires the LLM to invoke functions beyond its knowledge, which cannot be directly injected by SFT. Consequently, this leads to significant hallucination issues during inference. In contrast, CBR outperforms Naïve CBR, thanks to the strong foundational capabilities of LLMs. These capabilities are further enhanced by incorporating SFT, making CBR+SFT the strongest baseline. Among all the methods, the proposed CBR+Re4 achieves the highest FF1 score for both datasets, showing improvement of 3.30% and 6.80% over the best baseline, CBR+SFT, for DCN and SP, respectively. This validates the effectiveness of both the reranking based retrieval finetuning method and the reinforced reuse finetuning method.
@@ 4.3.1. Analyses for Reranking-based Retrieval Finetuning
As the key challenge for finetuning retriever models is lack of ground-truth labels, most advanced finetuning methods cannot be adapted to our setting. Thus, we analyze the proposed reranking-based retrieval finetuning method via an ablation study. The performance gap of CBR+Re4 without retrieval finetuning is reported for both datasets in Figure 5(a). The results show that the ablation causes a performance drop of approximately 1% in terms of FF1 score on both datasets. This highlights the effectiveness of the proposed reranking-based retrieval finetuning method.
@@ 4.4. Online Evaluation Results
We compare the previously deployed CBR+SFT system with two methods: (1) CBR+Re4 and (2) CBR+Re4 w/o Retain, an ablation variant of the CBR system where the Retain step is removed and the case bank remains fixed. It is worth noting that CBR+Re4 w/o Retain can be viewed as a variant of RAG, as it only reserves the Retrieve-Reuse process in CBR. Figure 7 depicts the average improvement in FF1 score as the number of online samples increases across two PDUs, DCN and SP. We observe that CBR+Re4 w/o Retain significantly underperforms CBR+Re4. This performance gap can be attributed to the dynamic nature of our scenario, where the software is continuously updated, and new modules are regularly introduced. Consequently, updating the case bank in an online manner is crucial for maintaining the effectiveness of test script generation. The Retain step in the CBR system addresses this need by enabling seamless integration of new information into the case bank. Furthermore, the proposed CBR+Re4 consistently outperforms the previously deployed CBR+SFT, achieving improvements in FF1 score of 2.75% in DCN and 10.08% in SP. These results validate the effectiveness of the proposed Re4 optimization method.
@@ 6. Conclusion
7. Limitation and Future Work
Firstly, we utilize bge-m3 with cosine similarity as the retriever in this work, which can be further improved by hybrid retrieval methods, multiple-stage retrieval methods, etc. Secondly, more comprehensive offline evaluation metrics may help us better benchmark the performance of the algorithms, such as function parameter accuracy, execution success rate, etc. In this work, we focus on the evaluation of function correctness, as this is the most important business metric based on our preliminary internal investigation. In contrast, execution success rate is a much more reliable evaluation metric; however, it is hard to calculate due to the complicated execution environment required by the target commercial software. Thus, there is still a long way to go towards fully automated functional testing, and this work aims to serve as a good starting point. Furthermore, the proposed CBR system can also be extended to broader decision-making scenarios in fields such as law, medicine, and finance. We plan to further explore the potentials in these domains in future work. Lastly, this paper focuses on function correctness and does not perform comprehensive user-centric evaluation, such as user acceptance rate, user satisfaction rate, time savings, etc. Investigation on these metrics may help us understand the bottleneck of the algorithms better.
Acknowledgements.
We truly thank the reviewers for their great effort in our submission. In this work, Siyuan Guo, Hechang Chen and Yi Chang are supported by National Natural Science Foundation of China through grants (624B2059), National Key R&D Program of China under Grant (No. 2023YFF0905400), National Natural Science Foundation of China through grants (U2341229, 61976102, U19A2065, 62476110), Key R&D Project of Jilin Province (No. 20240304200SF), and International Cooperation Project of Jilin Province (No. 20220402009GH).
@@ Appendix A Algorithmic Details
A.1. Business Metric Details
@@ Appendix B Why Does REINFORCE Work in Our RLFT Setting?
Now, we provide empirical evidence supporting the above claim. Since a smaller gradient variance corresponds to a smaller gradient norm, we follow prior work (Li et al., 2024) and plot the gradient norm of REINFORCE in our RLFT setting and RLHF setting in Figure 9. For RLHF, we follow (Li et al., 2024) to train the LLM with the consistent training setups as ours, where we adopt three prompt dataset (ultrafeedback, lmsys-chat-1m, sharegpt-en) and the reward model (UltraRM-13B) for RLHF. The results reveal that the gradient norm of REINFORCE in our RLFT settings remains below 1 across both datasets, whereas this value fluctuates between 10 and 4000 in RLHF. Thus, REINFORCE does not suffer from the high variance issue and can achieve the desired performance in our RLFT setting. Note that the gradient norm of REINFORCE in RLHF remains 0 in the later stage of training due to training collapse.
Figure 9. Comparison for the gradient norm of REINFORCE in RLFT and RLHF.
@@ Appendix C Discussion of the Compared RLHF Algorithms
In this section, we present a detailed discussion on how we tailor the state-of-the-art RLHF algorithms compared in the main body of the paper for our RLFT setting.
@@
where $T_{i}$ denotes the number of tokens in $i$-th generated script, $\hat{\rho}_{i,t}=\frac{\pi_{\theta}(\hat{y}_{i,t}|q,\mathcal{C}_{q},\hat{y}_{i,<t})}{\pi_{\theta_{\text{old}}}(\hat{y}_{i,t}|q,\mathcal{C}_{q},\hat{y}_{i,<t})}$ denotes the importance rate of the $t$-th token in $i$-th generated test script, $\hat{A}_{i,t}=\frac{r(\hat{y}_{i})-\text{mean}(\{r(\hat{y}_{1}),...,r(\hat{y}_{K})\})}{\text{std}(\{r(\hat{y}_{1}),...,r(\hat{y}_{K})\})}$ denotes the group-relative reward for advantage estimation, $\epsilon$ denotes the clipping parameter. GRPO achieves better performance than the proposed REINFORCE algorithm, benefiting from $K\times$ on-policy samples per query. Different from (Shao et al., 2024) that sets $K=64$, we set $K=4$ in our implementation to accommodate computational constraints, which brings 4×4$\times$4 × time and computational costs compared to REINFORCE.